# Disturbance Compensation for Safe Kinematic Control of Robotic Systems with Closed Architecture

### v1

@@ V Hardware Experiments
The inner-loop controller is configured as $\tau=-K_{p}({q}-{q}_{d})-K_{d}(\dot{q}-\dot{q}_{d})$, where $K_{p}=\mathrm{diag}([18,144,9])$ and $K_{d}=\mathrm{diag}([18,12,9])$. The gains are non-round values due to the actuator driver’s voltage-to-torque conversion ratio. The structure and true parameters remain unknown to our outer-loop controller. The nominal second-order system model (12) is used for the controller. The reference trajectory for joints $q_{1}$, $q_{2}$, and $q_{3}$ is $\begin{bmatrix}0.5\sin(t)&0.25\sin(2t)+0.534&0.25\sin(2t)+\frac{\pi}{2}\end{bmatrix}^{T}$. The tracking performance of the nominal inner-loop controller is illustrated in Fig. 4. Specifically, Fig. 4a depicts the trajectory tracking performance in the joint space, while Fig. 4b shows the trajectory tracking performance in the Cartesian space. The results indicate that the inner-loop controller exhibits poor tracking performance. We will demonstrate the improvement achieved by adding our approach to the tracking and safe control tasks.
We use the nominal model (12) throughout the experiments. A key parameter in the controller design is the control gain in (24), $b_{0}\triangleq\bar{M}^{-1}\bar{K}_{d}$, which depends on the nominal inertia matrix $\bar{M}$ and the inner-loop proportional gain. The nominal inertia $\bar{M}$ is adopted from [33]. Following a common ESO tuning rule, we tune each joint individually by initially setting $b_{0}$ large to ensure stability and then gradually reducing it to improve tracking accuracy. The final nominal control gain is $\mathrm{diag}([20,40,10])$. The observer and control bandwidths are set to $80~\mathrm{rad/s}$ and $10~\mathrm{rad/s}$, respectively, for each joint. We refer readers to [34] for additional instructions and theoretical support on tuning nominal control gain.
@@
V-A Tracking Control
We compare the tracking performance of ours, with [2]. The reason is that we share the same minimal assumptions about the inner-loop controller, which, to the best of our knowledge, are the most general in the literature.
V-A1 Scenario 1: Normal Case without Payload
The robot tracks a reference trajectory without a payload. The tracking performance of the learning-based adaptive controller [2] and ours is shown in Fig. 5. Fig. 5a compares the reference (dashed) and actual joint trajectories of $q_{1}$, $q_{2}$, and $q_{3}$, where blue and red lines represent our method and [2], respectively. The x-axis denotes time (s), and the y-axis denotes joint angles (rad). Fig. 5b shows the Cartesian trajectories in $x$, $y$, and $z$. The close alignment of the final and reference trajectories indicates that both controllers achieve accurate tracking despite model uncertainties and an imperfect inner loop.
The proposed method achieves tracking performance comparable to the learning-based adaptive controller in the no-payload experiment. As shown in Fig. 6, both approaches yield similarly smooth control signals. Transient responses occur initially, particularly in the second joint (Fig. 5a), due to the offset between the initial and reference positions. The learning-based controller exhibits a longer transient phase (see Figs. 5a and 6). This reflects a typical trade-off in adaptive control between transient duration and tracking accuracy.
@@ V-A1 Scenario 1: Normal Case without Payload
Figure 5: Comparative tracking precision analysis: (a) Root Mean squared error (RMSE) in joint space: Proposed ($4.1\times 10^{-3}$, $2.76\times 10^{-2}$, $7.6\times 10^{-3}$) radians vs. [2] ($4.3\times 10^{-3}$, $2.7\times 10^{-2}$, $5.7\times 10^{-3}$) radians for each joint. (b) Cartesian tracking error RMSE: Proposed ($2.3$ mm, $2.4$ mm, $15.5$ mm) vs. [2] ($2.9$ mm, $2.1$ mm, $16.1$ mm) for each joint. Data sampled at 1 kHz over a 10 s trajectory execution.
@@
V-A2 Scenario 2: Robustness Testing with Different Payloads
We evaluate controller robustness with payloads of 1 kg, 1.5 kg, and 2 kg attached to the end-effector. Each payload experiment is repeated five times for ANOVA analysis, with identical controller and observer parameters. Fig. 7a and Fig. 7b present the tracking error ANOVA results for both methods. As shown earlier, our approach achieves a shorter transient time than the learning-based adaptive controller. For clarity, transient data (first 5 s) are excluded, and the plots show steady-state tracking errors (after 5 s).
@@ V-A2 Scenario 2: Robustness Testing with Different Payloads
(b) ANOVA analysis for learning-based adaptive control [2]
@@
Figure 7: Steady-state tracking error comparison under different payloads. (a) ANOVA of ours. (b) ANOVA of the learning-based adaptive control [2]. The proposed method shows consistent tracking across payloads, as indicated by uniform box plots, while the adaptive method exhibits larger error variations, implying reduced robustness. Significance levels: ns ($p>0.05$), * ($p\leq 0.05$), ** ($p\leq 0.01$), *** ($p\leq 0.001$). Each experiment is repeated five times.
@@
As shown in Fig. 7a, the ANOVA results indicate that our method maintains statistically consistent errors across different payloads ($p>0.05$), demonstrating strong robustness, whereas the adaptive method shows significant variations with payload changes ($p<0.05$).
In summary, our approach offers several advantages over the learning-based adaptive control [2]. As shown in Fig. 5, it achieves comparable tracking performance without a payload, lower steady-state error, and stable performance under varying payloads (Fig. 7). It also exhibits a shorter transient response, as seen in the trajectories and control signals during the first second (Figs. 5 and 6). Moreover, our design is simpler than a neural network, requiring fewer tunable parameters and offering greater interpretability. Finally, the proposed disturbance compensation supports both tracking and safe control with state constraints, whereas [2] cannot handle constraints.
V-B Safe Control
It is worth noting that $\gamma$ in (53) critically influences the CBF’s behavior. A larger $\gamma$ drives the trajectory closer to the safety set boundary (more aggressive), while a smaller $\gamma$ keeps it farther away (more conservative). To achieve a safe yet efficient controller, we set $\gamma=10$ in all experiments.
We have conducted experiments in three scenarios: (1) Safety and Robustness Testing with Different Payloads, (2) Safety Testing Subject to External Disturbance, and (3) Safety Testing Subject to Gravity Disturbance. The first scenario will be described in detail, whereas the second and third scenarios are demonstrated in our accompanying video.
@@ V-B Safe Control
A virtual boundary is defined at $y_{0}=\begin{bmatrix}y_{1}&-0.1&y_{3}\end{bmatrix}^{T}$ as a spatial constraint for the end effector. It moves freely along the $x$ and $z$ axes but is constrained along the $y$ axis with a lower limit of 0.1-0.1 m. The system is tested with payloads of $1$, $1.5$, and $2$ kg attached to the end effector.
@@
Three approaches are compared: (1) conventional CBF with a nominal model (see Fig. 8); (2) our robust CBF with a nominal model and an ESO for disturbance estimation, compared to the DOB-CBF [16] (see Fig. 9a); and (3) our robust CBF with a nominal model, ESO for disturbance estimation, and estimation error bound, compared to the DOB-CBF with estimation error bound (see Fig. 9b).
@@
Fig. 8 shows the safety performance of the CBF using only the nominal second-order model (12), which neglects model errors, inner-loop uncertainty, and external disturbances. The safety function $h$ varies under different payloads, revealing limited robustness.
We implement the DOB-CBF [16] for comparison. The DOB-CBF estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$ in the CBF, while our method directly estimates $f$ in the original dynamics. Thus, the two observers target different quantities, and their gains are not directly comparable. For fairness, excluding error bounds, we tune both observers’ gains to achieve similar minimum $h$ magnitudes, indicating comparable safety performance. Fig. 9a compares the $h$ functions of the DOB-CBF and the proposed method, showing that:
@@
    A slight safety violation occurs due to disturbance estimation error, with minimum values of 0.597-0.597 mm for DOB-CBF and 0.366-0.366 mm for the proposed method, highlighting the need to account for estimation error bounds in both.
We collect experimental data for $\hat{b}_{e}(x,d)$ and $\hat{f}$ from the DOB and ESO, respectively. Their finite differences yield the maximum rates of change, used as estimation error bounds. Fig. 9b shows the resulting $h$ values after incorporating these bounds. Compared with Fig. 9a, we observe that:
@@
    The trajectories remain safe and consistent across different payloads, demonstrating the robustness of both methods.
@@
    Incorporating the estimation error bound raises $h$, keeping trajectories strictly within the safety set (positive $h$ values).
@@
    With error bounds, our method stays closer to the safety boundary than [16], indicating lower conservatism. This may be because [16] estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$, which includes a state-dependent Jacobian, making it more sensitive to measurement noise and compounding estimation errors.
VI Conclusion and future work

### v2

@@ VI Hardware Experiments
where $K_{p}=\mathrm{diag}([1,12,1])$ and $K_{d}=\mathrm{diag}([1,1,1])$. Again, this structure is unknown to our outer-loop controller. The nominal second-order system model (14) is used for the controller. The reference trajectory for joints $q_{1}$, $q_{2}$, and $q_{3}$ is $\begin{bmatrix}0.5\sin(t)&0.25\sin(2t)+0.534&0.25\sin(2t)+\frac{\pi}{2}\end{bmatrix}^{T}$. The tracking performance of the nominal inner-loop controller is illustrated in Fig. 5. Specifically, Fig. 5a depicts the trajectory tracking performance in the joint space, while Fig. 5b shows the trajectory tracking performance in the Cartesian space. The results indicate that the inner-loop controller exhibits poor tracking performance. We will demonstrate the improvement achieved by adding our approach to the tracking and safe control tasks.
The robot model and system parameters used here are adopted from our previous work [khalaf2019trajectory]. The nominal $\bar{K_{d}}$ is set to $\mathrm{diag}([1,1,1])$, and the nominal $\bar{B}$ is estimated through open-loop testing and fine-tuned as $\mathrm{diag}([20,40,10])$. The observer and control bandwidths are set to $80~\mathrm{rad/s}$ and $10~\mathrm{rad/s}$, respectively, for each joint.
@@
VI-A Robustness Testing under Different Nominal Gains
Although the default nominal $\bar{K}_{d}$ is set equal to the true value, we have conducted extensive tests with different $\bar{K}_{d}$ values to evaluate the robustness of our approach when the nominal value is inaccurate. Fig. 6 shows the tracking error sensitivity to $\bar{K}_{d}$, while keeping other parameters fixed. The results demonstrate remarkable gain robustness, with the tracking error remaining below 0.2 radians across a wide range of $0.6\times,10\times]$ the true $K_{d}$ value.
Figure 6: Comparison of tracking errors with different $K_{d}$ scaling factors using the nominal model (14). All results show a tracking error of less than 0.2 radians across a wide range of scaling factors: $0.6\times,10\times]$ the true $K_{d}$.
The determination strategy follows the tuning guidance discussed in Section IV-F. Initial stabilization is achieved by setting $b_{0}$ sufficiently high, followed by a gradual reduction to improve tracking accuracy, until instability is observed at low gain values (e.g., unstable performance at $0.5K_{d}$ observed in the experiments).
VI-B Tracking Control
In this section, we compare the tracking performance of our approach, using the control law designed in (26) with that of the state-of-the-art learning-based adaptive control subject to closed architecture [ahanda2022adaptive]. The reason for selecting [ahanda2022adaptive] for comparison is that we share the same minimal assumptions about the inner-loop controller, which, to the best of our knowledge, are the most general in the literature. In contrast, other works, such as [wang2019dynamic], assume more specific structures.
We demonstrate the comparisons in a normal scenario without a payload and another scenario with different payloads.
VI-B1 Scenario 1: Normal Case without Payload
In this scenario, the robot is required to follow the prescribed reference trajectory without a payload. The tracking performance of both the learning-based adaptive control [ahanda2022adaptive] and ours is shown in Fig. 7. The root mean squared errors (RMSEs) of the approaches under comparison are reported in Table II. Fig. 7a shows the joint reference and joint positions of three joints, $q_{1}$, $q_{2}$, and $q_{3}$. The dashed line is the reference trajectory, the blue line is the joint trajectory of ours, and the red line is the trajectory of [ahanda2022adaptive]. The x-axis represents time (in seconds), and the y-axis denotes the angular position of the joints (in radians). Fig. 7b illustrates the trajectory in the Cartesian space, with labeled axes $x$, $y$, and $z$. The close alignment of the final and reference trajectories suggests that both controllers can effectively track the desired trajectory despite model uncertainties and an imperfect inner-loop controller.
The tracking performance of ours is comparable to that of the learning-based adaptive controller in this experiment without a payload. As shown in Fig. 8, the control signal smoothness of the two approaches is also similar, with no significant difference. Transient states are observed at the beginning for both methods, especially for the second joint in Fig. 7a, due to the distance between the initial robot joint positions and the reference. The learning-based adaptive controller takes longer in the transient phase (see the state transition in Fig. 7a and the control variance in Fig. 8). We notice that the adaptive method always involves a trade-off between transient time and tracking accuracy; i.e., to achieve better tracking performance, it requires a longer transient time, and vice versa.
@@ VI-B1 Scenario 1: Normal Case without Payload
Figure 7: Comparative tracking precision analysis: (a) Root Mean squared error (RMSE) in joint space: Proposed ($4.1\times 10^{-3}$, $2.76\times 10^{-2}$, $7.6\times 10^{-3}$) radians vs. [ahanda2022adaptive] ($4.3\times 10^{-3}$, $2.7\times 10^{-2}$, $5.7\times 10^{-3}$) radians for each joint. (b) Cartesian tracking error RMSE: Proposed ($2.3$ mm, $2.4$ mm, $15.5$ mm) vs. [ahanda2022adaptive] ($2.9$ mm, $2.1$ mm, $16.1$ mm) for each joint. Data sampled at 1 kHz over a 10 s trajectory execution.
TABLE II: Comparison of RMSE in joint and Cartesian space
Joint 1 Joint 2 Joint 3
Root Mean Squared Error (RMSE) in Joint Space [rad]
Proposed $4.1\times 10^{-3}$ $2.76\times 10^{-2}$ $7.6\times 10^{-3}$
Proposed (zero-gravity) $1.3\times 10^{-3}$ $7.5\times 10^{-3}$ $2.4\times 10^{-3}$
Adaptive [ahanda2022adaptive] $4.3\times 10^{-3}$ $2.7\times 10^{-2}$ $5.7\times 10^{-3}$
x y z
Cartesian Tracking Error RMSE [mm]
Proposed $2.3$ $2.4$ $15.5$
Proposed (zero-gravity) $1.5$ $0.7$ $4.0$
Adaptive [ahanda2022adaptive] $2.9$ $2.1$ $16.1$
@@
Comparison with and without gravity compensation: As noted in Assumption1, the gravity compensation term is optional in the inner-loop controller, since in practice, robotic arms may operate in either zero-gravity mode (where gravity is already compensated by the inner-loop controller) or non-zero-gravity mode. Without loss of generality, we choose the non-zero-gravity mode setting; see (57) and (59) for the simulation and hardware experiments, respectively. We also evaluate a variant of our method with a zero-gravity inner-loop controller (referred to as “Proposed (zero-gravity))” in Table II). The results show that both controllers achieve superior performance, demonstrating the robustness and generality of our approach across two different configurations. The zero-gravity mode yields slightly better results, which is reasonable because zero-gravity typically implies less disturbance to compensate for, making the control task easier.
VI-B2 Scenario 2: Robustness Testing with Different Payloads
In this scenario, we test the robustness of different controllers with payloads of 1 kg, 1.5 kg, and 2 kg attached to the end-effector, shown in Fig. 9. Each experiment with a specific payload has been repeated five times for the ANOVA analysis. All controller and observer parameters have been kept the same. Fig. 10a and Fig. 10b show the ANOVA results of the tracking errors for the two approaches. As previously shown, our method exhibits a shorter transient time compared to the learning-based adaptive controller. For a more comprehensive comparison, the transient-phase data (first 5 seconds) were excluded, and the plots specifically present the steady-state tracking errors (from 5 seconds onward).
Figure 9: Experimental setup for robustness testing. Payloads of 1 kg (left), 1.5 kg (middle), and 2 kg (right) are attached to the end-effector.
@@ VI-B2 Scenario 2: Robustness Testing with Different Payloads
(b) ANOVA analysis for learning-based adaptive control [ahanda2022adaptive]
@@
Figure 10: Steady-state tracking error comparison with different payloads. (a) ANOVA analysis of the proposed method. (b) ANOVA analysis of the learning-based adaptive control [ahanda2022adaptive]. The proposed method maintains consistent tracking performance across different payloads, as evidenced by the uniform shape of the box plots; in contrast, the adaptive method shows significant variations in its error distribution, indicating compromised robustness. Significance levels: ns ($p>0.05$, not significant), * ($p\leq 0.05$), ** ($p\leq 0.01$), *** ($p\leq 0.001$). The experiments have been repeated five times.
@@
As shown in Fig. 10a, the significance testing from the ANOVA demonstrates that our method maintains statistically consistent error distributions across varying payload conditions ($p>0.05$), indicating strong robustness. In contrast, the adaptive methods exhibit significant variations due to payload changes ($p<0.05$).
In summary, our approach demonstrates significant advantages compared to the learning-based adaptive control [ahanda2022adaptive]:
  1. 1.
    As shown in Table II, the tracking performance of ours is comparable to that of the learning-based adaptive controller in this experiment with- out a payload.
  2. 2.
    Our approach exhibits lower steady-state error and achieves stable and consistent tracking performance under different payloads, as shown by the ANOVA test in Fig. 10.
  3. 3.
    Our approach has a shorter transient time, as shown in the trajectories and control signals during the first second in Fig. 7 and Fig. 8, respectively.
  4. 4.
    Our design is much simpler than a neural network, with far fewer parameters to tune and better interpretability, making it easier for users to perform troubleshooting.
  5. 5.
    Our disturbance compensation is not only used for tracking but can also be applied in safe control design with state constraints, whereas [ahanda2022adaptive] cannot handle any constraints.
However, we also acknowledge that [ahanda2022adaptive] unifies all components within a learning framework, whereas ours approach requires an additional step for estimating the nominal control gain.
VI-C Safe Control
In this section, we validate the safe control performance of our approach. It is essential to highlight that $\gamma$ in (56) is a crucial parameter related to the CBF’s behavior. A large $\gamma$ makes the trajectory approach closer to the safety set’s boundary, meaning a more aggressive behavior. A small $\gamma$ enforces the trajectory to stay far away from the safety set’s boundary, meaning a more conservative behavior. To design a safe and efficient controller that is desired in practice, we set $\gamma$ to a significantly high value of $10$ for all the experiments in this section.
As previously mentioned, disturbances can either help push the trajectory back into the interior of the safety set or cause it to escape the safety set. Both of these behaviors are undesirable. We will showcase these three scenarios in the following experiments: (1) Safety and Robustness Testing with Different Payloads, (2) Safety Testing Subject to External Disturbance, and (3) Safety Testing Subject to Gravity Disturbance.
VI-C1 Safety and Robustness Testing with Different Payloads
@@
A virtual boundary is set at $y_{0}=\begin{bmatrix}y_{1}&-0.1&y_{3}\end{bmatrix}^{T}$ as a spatial constraint for the end effector. The end effector is allowed unrestricted movement along the $x$ and $z$ axes but is constrained along the $y$ axis, with a lower limit set at 0.1-0.1 m. The system is evaluated with payloads of $1$ kg, $1.5$ kg, and $2$ kg attached to the end effector.
@@
We will present three approaches for comparison: (1) conventional CBF with a nominal model (see the safety function $h$ in Fig. 11); (2) our robust CBF with a nominal model and an ESO for disturbance estimation, comparared to the DOB-CBF [dacs2022robust] (see Fig. 12a); and (3) our robust CBF with a nominal model, ESO for disturbance estimation, and estimation error bound, compared to the DOB-CBF with estimation error bound (see Fig. 12b).
@@
Fig. 11 illustrates the safety performance of CBF with only a nominal second-order system model (14). This nominal model does not account for any uncertainty arising from model errors, an uncertain inner-loop controller, or external disturbances. The safety function $h$ exhibits variations due to external disturbances from varying payloads, indicating a lack of robust performance.
The details of implementing the DOB-CBF [dacs2022robust] have been included in Appendix B. The DOB-CBF estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$ in CBF, whereas our approach directly estimates $f$ in the original dynamic system. This leads to a fundamental difference: the two observers estimate different quantities, making direct comparison difficult. In particular, the key parameters—observer gains—cannot be directly compared because they pertain to two fundamentally different observers.
To ensure fairness, without including error bounds, we tune the observer gains for both approaches such that they yield similar minimum $h$ values (in terms of magnitude), suggesting comparable safety performance without error bounds. Fig. 12a illustrates the $h$ function comparison between the DOB-CBF and the proposed method. It is observed that:
@@
    Due to disturbance estimation error, a slight violation of the safety specification occurs, with the lowest point reaching 0.597-0.597 mm for DOB-CBF and 0.366-0.366 mm for the proposed method. This necessitates the consideration of estimation error bounds for both methods.
We collect experimental data for the estimated $\hat{b}_{e}(x,d)$ and $\hat{f}$ from the DOB and ESO, respectively. Their finite differences are computed to obtain the maximum values, which serve as bounds on their rates of change. Fig. 12b illustrates the resulting $h$ values after incorporating the estimation error bounds of the two approaches. Compared to the previous experiment shown in Fig. 12a, we have the following observations:
@@
    The trajectories are safe and consistent under different payloads, indicating robustness of both methods.
@@
    The incorporation of the estimation error bound elevates the value of $h$, making the trajectories always stay in the safety set; see the positive $h$ function values.
@@
    When incorporating error bounds, our approach remains closer to the safety boundary than that of [dacs2022robust], indicating reduced conservatism. A possible explanation is that [dacs2022robust] estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$, which involves a state-dependent Jacobian matrix. In a real robot with measurement noise, this introduces additional uncertainty, on top of the uncertainty from the disturbance $f$, potentially resulting in compounded estimation error. However, we acknowledge that the difference is not large, typically within the order of millimeter in $h$.
For practitioners using our approach, we offer the following recommendations regarding whether to include the error bound: First, if the CBF and disturbance compensation already yield satisfactory results—i.e., ensuring safety without being overly conservative—then incorporating the error bound may be unnecessary, particularly when the disturbance rate of change cannot be estimated within a tight range. Second, if performance is compromised due to estimation errors, then including the error bound is advised. However, it is crucial that the disturbance rate of change be carefully estimated; overly loose estimates can lead to over-conservatism.
VI-C2 Safety Testing Subject to External Disturbance
In this scenario, a constant 6V is applied to $q_{1}$, the waist joint, intentionally generating a continuous external perturbation that pushes the manipulator’s end-effector toward the virtual wall to evaluate the controller’s response. As shown in (58), the motor torque is controlled by an analog voltage.
Figure 13: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{-0.1}&\bm{y_{3}}\end{bmatrix}^{T}$. Trajectory comparison of CBF subject to external disturbance. The trajectory (blue) always stays in the safe zone. The unsafe trajectory (red) crosses the safety boundary.
Fig. 13 illustrates a trajectory comparison of CBFs with and without disturbance compensation. The red line shows the unsafe trajectory of the CBF that uses only the nominal model (14). While the tracking performance is satisfactory due to the nominal controller’s disturbance rejection capacity, the CBF is expected to intervene when the reference trajectory crosses the safety boundary. Unfortunately, the final trajectory still crosses the safety boundary because the CBF does not account for the external disturbance pushing the end-effector toward the unsafe region. In contrast, as shown in the blue line, our approach ensures high-performance tracking when safety is not a concern (see area $y=[0,0.3]$) and safe control, as the trajectory never crosses the safety boundary.
VI-C3 Safety Testing Subject to Gravity Disturbance
In this scenario, we set the virtual boundary as $y_{0}=\begin{bmatrix}y_{1}&y_{2}&-0.01\end{bmatrix}^{T}$. The end effector is allowed unrestricted motion along the $x$- and $y$-axes, but movement along the $z$-axis is constrained by a lower limit of 0.01-0.01 m. The disturbance caused by gravity will inevitably push the end effector toward the virtual boundary, leading to unsafe behavior. Remember that our nominal model (14), which has minimal model information, treats gravity as a disturbance.
Figure 14: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{y_{2}}&\bm{-0.01}\end{bmatrix}^{T}$. Trajectory comparison of CBFs subject to gravity disturbance. The safe trajectory (blue, ours) remains within the safe zone, while the unsafe trajectory (red, without disturbance compensation) crosses the safety boundary.
Fig. 14 shows the trajectory comparison of CBFs subject to gravity disturbance. The red line illustrates the unsafe trajectory of a CBF using only the nominal model (14). In contrast, our robust CBF significantly improves performance, as shown by the blue line, which consistently remains within the safe zone.
Computation complexity: The computational complexity is not the reason for removing joints 4, 5, and 6. Their removal is solely due to their redundancy in the 3D position control of the end-effector (i.e., the position of the wrist center). We break down the essential computations in our framework: (i) The observer in (24) involves only the simple integration of ordinary differential equations. It can easily scale to higher dimensions. (ii) The disturbance rejection controller in (26) has an explicit form without heavy computation. (iii) Our robust CBF in (56) is still based on standard convex Quadratic Programming (QP) solving, which is scalable to higher dimensions. The only additional terms are from the disturbance computed from the observer and the disturbance error bound (51). The disturbance error bound is computed offline. We emphasize that our hardware implementation operates at a 1 kHz frequency for both measurement sampling and control in dSPACE.
Possible extensions to other robotic systems: Our approach has the potential to be extended to other robotic systems, particularly commercial robots that face the same challenge addressed in this work: closed architectures with only kinematic controllers and inner-loop controllers implemented as unmodifiable firmware. The rationale behind this possible extension is that our dynamic system follows a general second-order Euler–Lagrange form, which is widely used in robotic motion control systems. In addition, the cascade control structure is common, with a kinematic controller in the outer loop and an inner-loop torque controller that typically includes a proportional term and an unknown function—potentially incorporating integral and dynamic compensation terms. However, we acknowledge that further customization may be necessary; for example, the presence of mismatched disturbances, which are common in many underactuated robots, poses additional challenges. Our recent work has demonstrated promising results in rejecting such mismatched disturbances in underactuated drones [chen2024quadrotor].
VII Conclusion and future work